Suppose that a cellular process resulted only in an increase in the entropy (S) of a system, but no change in enthalpy (H). What would the effect of this be on ΔG?
A) A change in S alone would not affect ΔG.
B) ΔG would be negative.
C) ΔG would be positive.
Answer #1

ΔG =ΔH-TΔS

T= temperature

ΔG= change in free energy

ΔH= change in enthalpy

ΔS= change in entropy

ΔS is posiitve, ( entropy has increased)

ΔH=0, (enthalpy did not change.)

ΔG =ΔH-TΔS

ΔG =0-TΔS

since  ΔS is posiitve, -TΔS is negative.

so the answer is B) ΔG would be negative.